# Reinforcement Twinning: from digital twins to model-based reinforcement learning

### v3

@@ Abstract
The concept of digital twins promises to revolutionize engineering by offering new avenues for optimization, control, and predictive maintenance. We propose a novel framework for simultaneously training the digital twin of an engineering system and an associated control agent. The training of the twin combines methods from adjoint-based data assimilation and system identification, while the training of the control agent combines model-based optimal control and model-free reinforcement learning. The training of the control agent is achieved by letting it evolve independently along two paths: one driven by a model-based optimal control and another driven by reinforcement learning. The virtual environment offered by the digital twin is used as a playground for confrontation and indirect interaction. This interaction occurs as an “expert demonstrator”, where the best policy is selected for the interaction with the real environment and “cloned” to the other if the independent training stagnates. We refer to this framework as Reinforcement Twinning (RT).
The framework is tested on three vastly different engineering systems and control tasks, namely (1) the control of a wind turbine subject to time-varying wind speed, (2) the trajectory control of flapping-wing micro air vehicles (FWMAVs) subject to wind gusts, and (3) the mitigation of thermal loads in the management of cryogenic storage tanks. The test cases are implemented using simplified models for which the ground truth on the closure law is available. The results show that the adjoint-based training of the digital twin is remarkably sample-efficient and completed within a few iterations. Concerning the control agent training, the results show that the model-based and the model-free control training benefit from the learning experience and the complementary learning approach of each other. The encouraging results open the path towards implementing the RT framework on real systems.
@@ 1 Introduction
Although the specific definition can vary considerably in different fields (Wagner et al., 2019; Barricelli et al., 2019; Chinesta et al., 2020; Rasheed et al., 2019; Ammar et al., 2022; Wright and Davidson, 2020; Tekinerdogan, 2022; van Beek et al., 2023; Haghshenas et al., 2023), a digital twin is generally seen as a virtual replica of a physical object (or system) capable of simulating its characteristic, functionalities and behaviour in real-time.
@@
Besides the obvious need for domain-specific knowledge and engineering, constructing such a replica involves various disciplines, from sensor technology and instrumentation for real-time monitoring to virtual/augmented reality and computer graphics for interfacing and visualization, and data-driven modelling, assimilation and machine learning for model updating and control. This work focuses on the last aspects and the required intersection of disciplines, with a particular emphasis on the need for real-time interaction. Real-time predictions are essential to deploying digital twins for control or monitoring purposes and require fast models. Fast models can be built from (1) macroscopic/lumped formulations, eventually enriched by data-driven ‘closure’ laws, (2) surrogate/template models, or (3) reduced-order models derived from high-fidelity simulations. In the context of system identifications for control purposes, the first would be referred to as ‘white’ models; the second would be referred to as ‘black box models’ while the third falls somewhere in between (see Schoukens and Ljung (2019) for the palette of grey shades in data-driven models). Regardless of the approach, the main conceptual difference between digital twinning and traditional engineering modelling is the continuous model update to synchronize the twin with the physical system and to provide a live representation of its state from noisy and partial observations.
@@
It is thus evident that the notion of digital twin lies at the intersection between established disciplines such as system identification (see Nelles (2001); Ljung (2008); Nicolao (2003)) and data assimilation (see Asch et al. (2016); Bocquet and Farchi (2023)). Both are significantly enhanced (and perhaps more intertwined) by the rapid growth and popularization of machine learning, as reviewed in the following section of this article. Both seek to combine observational data with a numerical model, and both involve an observation phase (known as training or calibration, depending on the field) in which the model is confronted with online data and a ‘prediction’ phase in which the model is interrogated. However, these disciplines are built around different problems and use different methods.
@@
Data assimilation is mostly concerned with the state estimation problem. It is assumed that the model is known (at least within quantifiable uncertainties), but observations are limited and uncertain. These uncertainties are particularly relevant because the system is chaotic and extremely high dimensional (e.g. atmospheric/oceanic models) and thus highly sensitive to the initial conditions from which predictions are requested. The traditional context is weather prediction and climate modelling (Lorenc, 1986; Wang et al., 2000; Pu and Kalnay, 2018; Lahoz et al., 2010), where the goal is to forecast the future evolution of a system from limited observations. The most common techniques are 4D variational assimilation (Talagrand and Courtier, 1987; Dimet et al., 2016; Ahmed et al., 2020) and, more recently, Ensemble Kalman methods (Evensen, 2009; Bocquet, 2011; Routray et al., 2016). Much effort is currently ongoing towards methods that combine these techniques (e.g. Kalnay et al. 2007; Lorenc et al. 2015.
@@
Nonlinear system identification is mostly concerned with the model identification problem. The system is subject to external inputs and observed through noisy measurements but is deterministic. The model can be fit into structures that are derived either from first principles or from general templates such as NRAX or Volterra models (Schoukens and Ljung, 2019), and the main goal is to predict how the system responds to actuation. The traditional context is control engineering, i.e. driving the system towards a desired trajectory. Besides the stronger focus on input-output relation (and thus the focus on actuated systems), the main difference with respect to traditional data assimilation is in the deterministic nature of the system to be inferred: errors in the initial conditions are not amplified as dramatically as, for example, in weather forecasting. Therefore, the resulting uncertainties remain within the bounds of model uncertainties and the state estimation is less critical than in data assimilation. Moreover, in many actuated systems, the forced response dominates over the natural one and the initial conditions are quickly forgotten. Tools for system identification are reviewed by Nelles (2001) and Suykens et al. (1996).
@@
Machine learning is progressively entering both communities with a wealth of general-purpose function approximations such as Artificial Neural Networks (ANNs, see Goodfellow et al. 2016) or Gaussian Processes (GPr, see Rasmussen and Williams 2005). More specifically, the fusion between these disciplines materializes through a combination of variational or ensemble tools from data assimilation to train dynamic models with model structures from machine learning to solve identification, forecasting or control problems. In this regard, the literature in control engineering has significantly anticipated the current developments (see neural controllers in Suykens et al. 1996; Norgaard et al. 2000), although their extension to broader literature has been limited by the mathematical challenges of nonlinear identification and the success of much simpler strategies based on piece-wise linearization or adaptive controllers (Astrom and Wittenmark, 1994; Sastry and Isidori, 1989).
@@
While machine learning enters the literature of data assimilation and system identification, cross-fertilization proceeds bilaterally. The need for identifying dynamic models from data is also growing in the literature on reinforcement learning, which is significantly drawing ideas from system identification and optimal control theory. Reinforcement learning is a subset of machine learning concerned with the training of an agent via trial and error to achieve a goal while acting on an environment (Sutton and Barto, 2018; Bertsekas, 2019). The framework differs from classic control theory because of its roots in sequential decision making: the environment to be controlled and the agent constitute a Markov Decision Process (MDP, Puterman (1994)) rather than a dynamical system with the usual state-space representation. Model-free approaches solely rely on input-output information and function approximations of the agent’s performance measures, while model-based approaches combine these with a model of the environment (Moerland et al., 2022a; Luo et al., 2022a). The model can be used as a forecasting tool, thus playing the same role as in model predictive control, or as a playground to accelerate the learning (Schwenzer et al., 2021). Model-free approaches have gained burgeoning popularity thanks to their success in video games (Szita, 2012; Mnih et al., 2013a) or natural language processing (Uc-Cetina et al., 2022), for which the definition of a system model is cumbersome. However, these algorithms have shown significant and somewhat surprising limitations for engineering problems governed by simple PDEs (Werner and Peitz, 2023; Pino et al., 2023).
@@ 2 Related Work
We briefly review recent developments in methods combining techniques from data assimilation (Section 2.1), system identification (Section 2.2) and model-based reinforcement learning (Section 2.3). We then briefly report on the main novelties of the proposed approach in Section 2.4.
@@ 2.1 Machine learning for Data Assimilation
An extensive overview of the state of the art of data assimilation (DA) is provided by Carrassi et al. (2017) while Cheng et al. (2023) give an overview of how machine learning (ML) is entering DA. Geer (2021) discuss the formal links between DA and ML in a Bayesian framework while Abarbanel et al. (2017) analyzes the link by building parallelism between time steps in a DA problem and layer labels in a deep ANN. As explored in Bocquet (2011), the formulation of hybrid DA-ML algorithms consists in introducing ML architectures either (1) to complement/correct the forecasting or the observation model or (2) to replace at least one of them. An early attempt at DA-ML hybridization within the first class was proposed by Tang and Hsieh (2001), who used a variational approach to train an ANN that replaces missing dynamical equations. More recent approaches within the first category are proposed by Arcucci et al. (2021) while Buizza et al. (2022) presents several approaches in both categories. In particular, Arcucci et al. (2021) implements a Recurrent Neural Network (RNN,Madhavan 1993) to correct the forecasting model, while Buizza et al. (2022) use convolutional neural networks (CNNs) to improve the observations passed to a Kalman filter. Within the second class of hybrid DA-ML methods, Brajard et al. (2020) and Buizza et al. (2022) use a Kalman filter to train the parameters of an ANN that acts as a forecasting model. Outside geophysical sciences, the use of ANNs as a forecasting tool in assimilation frameworks is spreading in economics (Khandelwal et al., 2021) and epidemiology (Nadler et al., 2020). Finally, within the literature on data assimilation, several works have combined ML and DA strategies for model identification (Bocquet et al., 2019; Ayed et al., 2019).
@@ 2.2 Machine learning for System Identification and Control
Recent reviews on deep learning for system identifications are proposed by Ljung et al. (2020) and Pillonetto et al. (2023) while earlier reviews are provided by Ljung (2008), Nicolao (2003). System identification has been historically influenced by developments in machine learning and adopted the use of ANNs for system modelling already in the 90’s (Chen et al., 1990; Zhang and Moore, 1991; Sjöberg et al., 1994). Early works (see also Suykens et al. 1996 and Norgaard et al. 2000) focused on the use of feedforward neural networks (FNNs) or recurrent neural networks (RNNs); these can be seen as special variants of nonlinear autoregressive and nonlinear state-space models. The popularization of ML has significantly enlarged the zoology of ANNs architectures used as model structures. The most popular examples include variants of the RNN, such as long-short-term memory (LSTM) networks (Hochreiter and Schmidhuber, 1997) and Echo State Networks (ESN, Jaeger and Haas (2004)), as well as convolutional neural networks (CNNs,LeCun et al. (1989)) and the latest developments in neural ODEs (also known as ODE-net,Chen et al. (2018)).
Important contributions within the first class of approaches are the works by Ljung et al. (2020), Gonzalez and Yu (2018) and Bucci et al. (2018) who use LSTM networks (Hochreiter and Schmidhuber, 1997). These are variants of RNNs, which are very popular in natural language processing and use three gates (forget/input/output) to preserve information over longer time steps, thus allowing for better handling of long-term dependencies. Canaday et al. (2020) used ESNs, which gained popularity for forecasting in stochastic systems. These are variants of RNNs implementing the reservoir computing formalism, i.e. ANNs in which weights and biases are selected randomly. Andersson et al. (2019) used CNNs, which gained popularity in image recognition and allow for identifying patterns at different scales, while Ayed et al. (2019) and Rahman et al. (2022) use ODE-nets, recently proposed as a new paradigm for neural state-space modelling. ODE-nets treat the layers of an ANN as intermediate time steps of a multistep ODE solver and are commonly trained using the adjoint method, a ubiquitous tool in data assimilation. An overview of the challenges of learning models of dynamical systems from data has been recently presented by Bucci et al. (2023), who also propose a “gradual” approach which seeks to build models of gradually increasing complexity.
@@ 2.3 Model Based Reinforcement Learning and Optimal Control
Recent reviews of model-based reinforcement learning (MBRL) are proposed by Chatzilygeroudis et al. (2019), Luo et al. (2022b) and Moerland et al. (2022b). MBRL uses a model of the environment to guide the agent, with the goal of increasing its sample efficiency (i.e. reducing the number of trials required to learn a task). A model allows the agent to sample the environment with arbitrary state-action values. In contrast, a model-free approach can only sample visited states, eventually stored in a replay buffer to allow multiple re-uses (Haarnoja et al., 2018; Lillicrap et al., 2019). In the reinforcement learning terminology, a model allows the agent to plan, giving the possibility to imagine the consequences of actions and, more generally, update policy and/or state-action value predictors without interacting with the environment.
@@
MBRL methods can be classified based on how the model is built and how it is used. Concerning the approaches for model constructions, a large arsenal of function approximations has been deployed in the literature. ANNs (Hunt et al., 1992; Kurutach et al., 2018; Nagabandi et al., 2017), Gaussian Processes (GPr,Deisenroth and Rasmussen (2011); Boedecker et al. (2014a)) and time-varying linear models (Boedecker et al., 2014b) are some examples. Concerning model usage, the simplest approach consists of using it solely to sample the state-action space, while more advanced approaches use the model in the training steps. One of the seminal MBRL approaches is the Dyna algorithm proposed by Sutton (1991). This uses experience (i.e. trajectories in the state-action space and associated rewards) collected from both environment and model. More sophisticated approaches use the model to “look ahead”, i.e. predicting the future evolution of the system under a given policy. A notable example is the Monte Carlo Tree Search in the AlphaGO (Silver et al., 2016, 2018) algorithm that reached superhuman performances in the game of Go. Moving from Markov Decision Processes to dynamical systems, the boundaries between MBRL and Model Predictive Control (MPC) become blurred, particularly when the latter uses black-box models (see Hedengren et al. 2014; Raw 2000 and Nagabandi et al. 2017 or Weber et al. 2017).
@@
Several authors have proposed methods to leverage physical insights in the models for MBRL, using physics-informed neural networks (Raissi et al., 2019) as Liu and Wang (2021) or Deep Lagrangian Networks (DeLaN Lutter et al. 2019) as Ramesh and Ravindran (2023). A more extreme approach is proposed by Lutter et al. (2020), who uses ‘white-box’ models (eventually enhanced as proposed by Lutter et al. 2020) in MBRL. This work shows a traditional physics-based model trained via machine learning methods. The opposite case is provided by Liu and MacArt (2023), who uses an ANN agent trained with adjoint-based optimization. An overview of methods to integrate physics-based models and machine learning is provided by Baker et al. (2019) and Willard et al. (2020). Regardless of its structure, a differentiable model of the environment can be trained using optimal control theory (Stengel, 1994).

### v4

@@ Abstract
The concept of digital twins promises to revolutionize engineering by offering new avenues for optimization, control, and predictive maintenance. We propose a novel framework for simultaneously training the digital twin of an engineering system and an associated control agent. The training of the twin combines methods from adjoint-based data assimilation and system identification, while the training of the control agent combines model-based optimal control and model-free reinforcement learning. The training of the control agent is achieved by letting it evolve independently along two paths: one driven by a model-based optimal control and another driven by reinforcement learning. The virtual environment offered by the digital twin is used as a playground for confrontation and indirect interaction. This interaction occurs as an “expert demonstrator”, where the best policy is selected for the interaction with the real environment and “cloned” to the other if the independent training stagnates. We refer to this framework as Reinforcement Twinning (RT). The framework is tested on three vastly different engineering systems and control tasks, namely (1) the control of a wind turbine subject to time-varying wind speed, (2) the trajectory control of flapping-wing micro air vehicles (FWMAVs) subject to wind gusts, and (3) the mitigation of thermal loads in the management of cryogenic storage tanks. The test cases are implemented using simplified models for which the ground truth on the closure law is available. The results show that the adjoint-based training of the digital twin is remarkably sample-efficient and completed within a few iterations. Concerning the control agent training, the results show that the model-based and the model-free control training benefit from the learning experience and the complementary learning approach of each other. The encouraging results open the path towards implementing the RT framework on real systems.
@@
[label1]organization=von Karman Institute,city=Rhode-St-Genése, postcode=1640, country=Belgium
@@
[label2]organization=Vrije Universiteit Brussel (VUB), Department of Mechanical Engineering,city=Elsene, Brussels, postcode=1050, country=Belgium
[…]
[label3]organization=Université Libre de Bruxelles,addressline=Av. Franklin Roosevelt 50, city=Brussels, postcode=1050, country=Belgium
[…]
[label4]organization=University of Ghent,addressline=Sint-Pietersnieuwstraat 41, city=Ghent, postcode=9000, country=Belgium
[…]
[label5]organization=Institute of Mechanics, Materials, and Civil Engineering (iMMC), Université Catholique de Louvain,addressline=, city=Louvain-la-Neuve, postcode=1348, country=Belgium
@@ 1 Introduction
Although the specific definition can vary considerably in different fields [Wagner et al., 2019, Barricelli et al., 2019, Chinesta et al., 2020, Rasheed et al., 2019, Ammar et al., 2022, Wright and Davidson, 2020, Tekinerdogan, 2022, van Beek et al., 2023, Haghshenas et al., 2023], a digital twin is generally seen as a virtual replica of a physical object (or system) capable of simulating its characteristic, functionalities and behaviour in real-time.
@@
Besides the obvious need for domain-specific knowledge and engineering, constructing such a replica involves various disciplines, from sensor technology and instrumentation for real-time monitoring to virtual/augmented reality and computer graphics for interfacing and visualization, and data-driven modelling, assimilation and machine learning for model updating and control. This work focuses on the last aspects and the required intersection of disciplines, with a particular emphasis on the need for real-time interaction. Real-time predictions are essential to deploying digital twins for control or monitoring purposes and require fast models. Fast models can be built from (1) macroscopic/lumped formulations, eventually enriched by data-driven ‘closure’ laws, (2) surrogate/template models, or (3) reduced-order models derived from high-fidelity simulations. In the context of system identifications for control purposes, the first would be referred to as ‘white’ models; the second would be referred to as ‘black box models’ while the third falls somewhere in between (see Schoukens and Ljung [2019] for the palette of grey shades in data-driven models). Regardless of the approach, the main conceptual difference between digital twinning and traditional engineering modelling is the continuous model update to synchronize the twin with the physical system and to provide a live representation of its state from noisy and partial observations.
@@
It is thus evident that the notion of digital twin lies at the intersection between established disciplines such as system identification (see Nelles [2001], Ljung [2008], Nicolao [2003]) and data assimilation (see Asch et al. [2016], Bocquet and Farchi [2023]). Both are significantly enhanced (and perhaps more intertwined) by the rapid growth and popularization of machine learning, as reviewed in the following section of this article. Both seek to combine observational data with a numerical model, and both involve an observation phase (known as training or calibration, depending on the field) in which the model is confronted with online data and a ‘prediction’ phase in which the model is interrogated. However, these disciplines are built around different problems and use different methods.
@@
Data assimilation is mostly concerned with the state estimation problem. It is assumed that the model is known (at least within quantifiable uncertainties), but observations are limited and uncertain. These uncertainties are particularly relevant because the system is chaotic and extremely high dimensional (e.g. atmospheric/oceanic models) and thus highly sensitive to the initial conditions from which predictions are requested. The traditional context is weather prediction and climate modelling [Lorenc, 1986, Wang et al., 2000, Pu and Kalnay, 2018, Lahoz et al., 2010], where the goal is to forecast the future evolution of a system from limited observations. The most common techniques are 4D variational assimilation [Talagrand and Courtier, 1987, Dimet et al., 2016, Ahmed et al., 2020] and, more recently, Ensemble Kalman methods [Evensen, 2009, Bocquet, 2011, Routray et al., 2016]. Much effort is currently ongoing towards methods that combine these techniques (e.g. Kalnay et al. 2007, Lorenc et al. 2015.
@@
Nonlinear system identification is mostly concerned with the model identification problem. The system is subject to external inputs and observed through noisy measurements but is deterministic. The model can be fit into structures that are derived either from first principles or from general templates such as Nonlinear AutoRegressive models with exogenous inputs (NARX, Lin et al. [1996]) or Volterra models [Schoukens and Ljung, 2019], and the main goal is to predict how the system responds to actuation. The traditional context is control engineering, i.e. driving the system towards a desired trajectory. Besides the stronger focus on input-output relation (and thus the focus on actuated systems), the main difference with respect to traditional data assimilation is in the deterministic nature of the system to be inferred: errors in the initial conditions are not amplified as dramatically as, for example, in weather forecasting. Therefore, the resulting uncertainties remain within the bounds of model uncertainties and the state estimation is less critical than in data assimilation. Moreover, in many actuated systems, the forced response dominates over the natural one and the initial conditions are quickly forgotten. Tools for system identification are reviewed by Nelles [2001] and Suykens et al. [1996].
@@
Machine learning is progressively entering both communities with a wealth of general-purpose function approximations such as Artificial Neural Networks (ANNs, see Goodfellow et al. 2016) or Gaussian Processes (GPr, see Rasmussen and Williams 2005). More specifically, the fusion between these disciplines materializes through a combination of variational or ensemble tools from data assimilation to train dynamic models with model structures from machine learning to solve identification, forecasting or control problems. In this regard, the literature in control engineering has significantly anticipated the current developments (see neural controllers in Suykens et al. 1996, Norgaard et al. 2000), although their extension to broader literature has been limited by the mathematical challenges of nonlinear identification and the success of much simpler strategies based on piece-wise linearization or adaptive controllers [Astrom and Wittenmark, 1994, Sastry and Isidori, 1989].
@@
While machine learning enters the literature of data assimilation and system identification, cross-fertilization proceeds bilaterally. The need for identifying dynamic models from data is also growing in the literature on reinforcement learning, which is significantly drawing ideas from system identification and optimal control theory. Reinforcement learning is a subset of machine learning concerned with the training of an agent via trial and error to achieve a goal while acting on an environment [Sutton and Barto, 2018, Bertsekas, 2019]. The framework differs from classic control theory because of its roots in sequential decision making: the environment to be controlled and the agent constitute a Markov Decision Process (MDP, Puterman [1994]) rather than a dynamical system with the usual state-space representation. Model-free approaches solely rely on input-output information and function approximations of the agent’s performance measures, while model-based approaches combine these with a model of the environment [Moerland et al., 2022a, Luo et al., 2022a]. The model can be used as a forecasting tool, thus playing the same role as in model predictive control, or as a playground to accelerate the learning [Schwenzer et al., 2021]. Model-free approaches have gained burgeoning popularity thanks to their success in video games [Szita, 2012, Mnih et al., 2013a] or natural language processing [Uc-Cetina et al., 2022], for which the definition of a system model is cumbersome. However, these algorithms have shown significant and somewhat surprising limitations for engineering problems governed by simple PDEs [Werner and Peitz, 2023, Pino et al., 2023].
@@ 2 Related Work
This section locates the contribution of this work in the broad literature of the intersected fields. We briefly review recent developments in methods combining machine learning with data assimilation (Section 2.1) and system identification (Section 2.2). We then move towards the cross-fertilization in the context of model-based reinforcement learning (Section 2.3) and focus on recent developments on the combination between model-based and model-free reinforcement learning (Section 2.4). Finally, Section 2.5 briefly reports on the main novelties of the proposed approach.
@@ 2.1 Machine Learning for Data Assimilation
An extensive overview of the state of the art of data assimilation (DA) is provided by Carrassi et al. [2017] while Cheng et al. [2023] give an overview of how machine learning (ML) is entering DA. Geer [2021] discuss the formal links between DA and ML in a Bayesian framework while Abarbanel et al. [2017] analyzes the link by building parallelism between time steps in a DA problem and layer labels in a deep ANN. As explored in Bocquet [2011], the formulation of hybrid DA-ML algorithms consists in introducing ML architectures either (1) to complement/correct the forecasting or the observation model or (2) to replace at least one of them. An early attempt at DA-ML hybridization within the first class was proposed by Tang and Hsieh [2001], who used a variational approach to train an ANN that replaces missing dynamical equations. More recent approaches within the first category are proposed by Arcucci et al. [2021] while Buizza et al. [2022] presents several approaches in both categories. In particular, Arcucci et al. [2021] implements a Recurrent Neural Network (RNN,Madhavan 1993) to correct the forecasting model, while Buizza et al. [2022] use convolutional neural networks (CNNs) to improve the observations passed to a Kalman filter. Within the second class of hybrid DA-ML methods, Brajard et al. [2020] and Buizza et al. [2022] use a Kalman filter to train the parameters of an ANN that acts as a forecasting model. Outside geophysical sciences, the use of ANNs as a forecasting tool in assimilation frameworks is spreading in economics [Khandelwal et al., 2021] and epidemiology [Nadler et al., 2020]. Finally, within the literature on data assimilation, several works have combined ML and DA strategies for model identification [Bocquet et al., 2019, Ayed et al., 2019].
@@ 2.2 Machine learning for System Identification and Control
Recent reviews on deep learning for system identifications are proposed by Ljung et al. [2020] and Pillonetto et al. [2023] while earlier reviews are provided by Ljung [2008], Nicolao [2003]. System identification has been historically influenced by developments in machine learning and adopted the use of ANNs for system modelling already in the 90’s [Chen et al., 1990, Zhang and Moore, 1991, Sjöberg et al., 1994]. Early works (see also Suykens et al. 1996 and Norgaard et al. 2000) focused on the use of feedforward neural networks (FNNs) or recurrent neural networks (RNNs); these can be seen as special variants of nonlinear autoregressive and nonlinear state-space models. The popularization of ML has significantly enlarged the zoology of ANNs architectures used as model structures for state-spacem odeling. The most popular examples include variants of the RNN, such as long-short-term memory (LSTM) networks [Hochreiter and Schmidhuber, 1997] and Echo State Networks (ESN, Jaeger and Haas [2004]), as well as convolutional neural networks (CNNs,LeCun et al. [1989]) and the latest developments in neural ODEs (also known as ODE-net,Chen et al. [2018]).
Important contributions within the first class of approaches are the works by Ljung et al. [2020], Gonzalez and Yu [2018] and Bucci et al. [2018] who use LSTM networks [Hochreiter and Schmidhuber, 1997]. These are variants of RNNs, which are very popular in natural language processing and use three gates (forget/input/output) to preserve information over longer time steps, thus allowing for better handling of long-term dependencies. Canaday et al. [2020] used ESNs, which gained popularity for forecasting in stochastic systems. These are variants of RNNs implementing the reservoir computing formalism, i.e. ANNs in which weights and biases are selected randomly. Andersson et al. [2019] used CNNs, which gained popularity in image recognition and allow for identifying patterns at different scales, while Ayed et al. [2019] and Rahman et al. [2022] use ODE-nets, recently proposed as a new paradigm for neural state-space modelling. ODE-nets treat the layers of an ANN as intermediate time steps of a multistep ODE solver and are commonly trained using the adjoint method, a ubiquitous tool in data assimilation. An overview of the challenges of learning models of dynamical systems from data has been recently presented by Bucci et al. [2023], who also propose a “gradual” approach which seeks to build models of gradually increasing complexity.
@@ 2.3 Model Based Reinforcement Learning and Optimal Control
Recent reviews of model-based reinforcement learning (MBRL) are proposed by Chatzilygeroudis et al. [2019], Luo et al. [2022b] and Moerland et al. [2022b]. MBRL uses a model of the environment to guide the agent, with the goal of increasing its sample efficiency (i.e. reducing the number of trials required to learn a task). A model allows the agent to sample the environment with arbitrary state-action values. In contrast, a model-free approach can only sample visited states, eventually stored in a replay buffer to allow multiple re-uses [Haarnoja et al., 2018, Lillicrap et al., 2019]. In the reinforcement learning terminology, a model allows the agent to plan, giving the possibility to imagine the consequences of actions and, more generally, update policy and/or state-action value predictors without interacting with the environment.
@@
MBRL methods can be classified based on how the model is built and how it is used. Concerning the approaches for model constructions, a large arsenal of function approximations has been deployed in the literature. ANNs [Hunt et al., 1992, Kurutach et al., 2018, Nagabandi et al., 2017], Gaussian Processes (GPr,Deisenroth and Rasmussen [2011], Boedecker et al. [2014a]), time-varying linear models [Boedecker et al., 2014b] are some examples. Concerning model usage, the simplest approach consists of using it solely to sample the state-action space, while more advanced approaches use the model in the training steps [Janner et al., 2021, Gu et al., 2016]. One of the seminal MBRL approaches is the Dyna algorithm proposed by Sutton [1991]. This uses experience (i.e. trajectories in the state-action space and associated rewards) collected from both environment and model. More sophisticated approaches use the model to “look ahead”, i.e. predicting the future evolution of the system under a given policy. A notable example is the Monte Carlo Tree Search in the AlphaGO [Silver et al., 2016, 2018] algorithm that reached superhuman performances in the game of Go. Moving from Markov Decision Processes to dynamical systems, the boundaries between MBRL and Model Predictive Control (MPC) become blurred, particularly when the latter uses black-box models (see Hedengren et al. 2014, Raw 2000 and Nagabandi et al. 2017 or Weber et al. 2017).
@@
Several authors have proposed methods to leverage physical insights in the models for MBRL, using physics-informed neural networks [Raissi et al., 2019] as Liu and Wang [2021] or Deep Lagrangian Networks (DeLaN Lutter et al. 2019) as Ramesh and Ravindran [2023]. A more extreme approach is proposed by Lutter et al. [2020], who uses ‘white-box’ models (eventually enhanced as proposed by Lutter et al. 2020) in MBRL. This work shows a traditional physics-based model trained via machine learning methods. The opposite case is provided by Liu and MacArt [2023], who uses an ANN agent trained with adjoint-based optimization. An overview of methods to integrate physics-based models and machine learning is provided by Baker et al. [2019] and Willard et al. [2020]. Regardless of its structure, a differentiable model of the environment can be trained using optimal control theory [Stengel, 1994].
2.4 The Hybridization of Model-Based and Model-Free Control
The idea of combining model-based and model-free control methods is gaining momentum in the literature of robotics and control and is generally referred to as hybrid learning [Pinosky et al., 2023]. Hybrid learning seeks to combine the model-based methods’ sample efficiency with the model-free methods’ robustness (see Dulac-Arnold et al. [2019]). Hybrid approaches differ in how the model-based and model-free policies are chosen or merged. An early idea proposed by Yamada et al. [1997] consisted of keeping these separate and independent, with procedures known as Hierarchical Mixtures of Experts (HME) [Jordan and Jacobs, 1994, Tham, 1995] used to identify the best of the two at a given state. More recently, a stronger interaction was proposed by Freed et al. [2024], who used a model-based policy to bound a model-free policy, with the goal of guaranteeing the stability of the controller even during the exploration. A similar approach is developed in the context of continuous reinforcement learning [Çağatay Yıldız et al., 2021, Lutter et al., 2021, Doya, 2000] by Pinosky et al. [2023]. Alternatively, if models are cast in a locally linear or linear-quadratic form, iterative optimal control methods [Tassa et al., , Van Den Berg, ] have been used to generate a pool of “expert” local policies from which a model-free learner must generalize and improve upon [Levine and Koltun, 2013, Levine and Abbeel, 2014, Chebotar et al., 2017, Kahn et al., 2017], or to warm-start a model-free policy search step [Qu et al., 2020]. [Abbeel et al., 2006] proposed an approach decoupling the policy evaluation and policy improvement steps, thereby developing a hybrid method in which one step is carried out in a model-free approach and the other proceeds in a model-based approach.
An alternative path is to obtain a hybridization in the value function space. Value functions aim to predict each state’s value and quantify how well a controller can perform from a given state of the system; value function approximators can thus be used to guide optimal controllers as proposed by [Pong et al., 2020]. Models can also be used to generate roll-outs to enhance model-free value estimation methods [Heess et al., 2015], potentially weighting the collected data according to the trust in the model at hand [Feinberg et al., 2018]. Other combinations propose the use of control-theoretic tools to provide guarantees on model-free methods. An example can be found in Berkenkamp et al. [2017], in which results on Lyapunov stability theory are used to identify safe regions in the state-space [Richards et al., 2018] in which the controller is constrained to operate. In addition, if considerations on robustness are also made, this effectively leads to an $H_{\infty}$ model-free RL approach [Han et al., ].